The patient weighs 700 grams. You are given orders to administer a medication 1 mEq/kg IV. The medication is available in 0.5 mEq/mL. The dose to be given would be _______mL. (record your answer using one decimal place)

if possible...please use the "divided over have" method because thats the one I use and it will help me better understand
Answer #1

The patient weighs 700 grams. =0.7 kg (As 1 kg= 1000 gm)

Ordered medication= 1 mEq/kg IV.

As per patient's weight required medication= for 1 kg weight reqired= 1 mEq

for 0.7 kg required = 0.7 mEq IV

Available =0.5 mEq/mL.

The dose to be given would be _______mL.

For 0.5 mEq required ml= 1

for 1 mEq required ml= 1/0.5

For 0.7 mEq required ml= (1/0.5) X 0.7= 1.4 ml

Hence 1.4 ml of the available medication should be deliverd by IV route to fulfill the order